Right triangle RST has vertices R(-2,-1) and S(0,1). If the measure of angleR is 90 degrees, which is teh equation of lineRT?
A. y = -5x + 2
B. y = -x - 3
C. y = -x-1
D. y = 5x - 3

Answer by oscargut(2103)

RS has equaton y=x+1, then RT has 1st coefficient -1 then RT: y=-x+b
but R is in RT then -2=-(-1)+b then b=-3 then RT: y=-x-3
option B is the correct